OpenStudy (anonymous):

Write 14% as a fraction and as a decimal.

OpenStudy (anonymous):

0.14

OpenStudy (anonymous):

14/100 = 7/50. 7/50 = 0.14

OpenStudy (anonymous):

14 /100 = 7/50 and .14

OpenStudy (anonymous):

hope it helps